OpenStudy (anonymous):

What is the x and y-intercepts? f(x)=((x-1)(x+9))/((-x-4)(9x-1))

OpenStudy (anonymous):

y intercept = f(0) x intercepts = f(x)=0

OpenStudy (anonymous):

y int = -9/4

OpenStudy (anonymous):

x intercepts will be when your numerator = 0 so -9,1

OpenStudy (anonymous):

Thanks so much Outkast3r09!!

OpenStudy (anonymous):

and when you have an asymptote i believe too so you should do the denominator also i believe

OpenStudy (anonymous):

so all together you should get x intercepts at -9,1,1/9,-4

OpenStudy (anonymous):

where 1/9 and -4 are retricemptotes

OpenStudy (anonymous):

asymptotes lol

OpenStudy (anonymous):

thanks bro!